OpenStudy (anonymous):

Write the equation of the line that is perpendicular to the line 2x − 3y = 3 and passes through the point (−8, 2).

OpenStudy (anonymous):

Change your equation to standard form: y=mx +b ------>3y=2x-3 y=2/3x-1 When (L1) perpencular with the line you have therefore its m=-3/2 (L1): y= -3/2x+b Passes(-8,2): 2=(-3/2)(-8)+b 2= 12+b b= 2-12=-10 (L1): y=-3/2x -10